OpenStudy (anonymous):

what about Q1,Q2,Q3. does somebody got anything??

OpenStudy (anonymous):

Q4 p1=V*R2/(R1+R2)*(1-e^(-(R1+R2)*t/(R1*R2*C))) p2=(V*R2/(R1+R2))*cos(1/sqrt(L*C)*(t-T)) p3=C*(V*R2/(R1+R2))*sin((t-T)*1/sqrt(L*C))/sqrt(L*C)

OpenStudy (anonymous):

Q5 E V B Y F Z

OpenStudy (anonymous):

Q6 with the values :V0=10V, T=1ms, Rw=2Ω, Lw=1mH and RL=3Ω 0.2 0.2 ? ?

OpenStudy (anonymous):

give best response if u get a green....

OpenStudy (anonymous):

For q1,q2 ,check out closed questions. question 3 is nothing but homework 8 H8P1 IMPULSE

OpenStudy (anonymous):

ok thanks.. if i get a green i ll post them here.. to be all together..

OpenStudy (anonymous):

Q2 p1: Vs as vin<vt p2: v0 = vs-kr(e^((vin-vt)/vn)-1) p3: vin = ln(vs/2kr +1)vn+vt p4: vin=sqrt(vs/kr)+vt p5: -kre^((vin-vt)/vn) p6: -kr(vin-vt)

OpenStudy (anonymous):

Q1 a)ia=2*i1*R/(2*R+rm) and vo=ia*rm b)iA =(-i2)*R/(2*R+rm) and V0=ia*rm c) part a) + part b) d)2R

OpenStudy (anonymous):

not a best response so far??????? i think thats not fair..

OpenStudy (anonymous):

b)iA =(-i2)*R/(2*R+rm) and V0=ia*rm --> -2*i or -i^2

OpenStudy (anonymous):

its in2 from the values its been given

OpenStudy (anonymous):

GIVE ME THE FORMULas 4 th question

OpenStudy (anonymous):

I'm getting error for Q2 p2 please help.

OpenStudy (anonymous):

all formullas posted...... http://openstudy.com/study?source=email#/updates/50d4a24de4b0d6c1d5417e42